OpenStudy (anonymous):

int {[sinˆ2(x)]/[sin^4(x)+1]}*sin(2x)dx

OpenStudy (anonymous):

???

OpenStudy (dumbcow):

sin(2x) = 2sinx cosx \[= 2\int\limits_{}^{}\frac{\sin^{3} x \cos x}{\sin^{4} x +1}\] then use substitution u = sin^4 du = 4sin^3 cos

OpenStudy (anonymous):

thank you so much

OpenStudy (dumbcow):

no problem:)